OpenStudy (anonymous):

Let f(x) = 8+sqrt(5 x-5) . Find f^(-1)(x)

OpenStudy (anonymous):

y = 8 + sqrt(5x-5) x = 8 + sqrt(5y-5) x = 13sqrt(5y) x = 13^2 * 5y^2 x = 65x^4 Can someone check my steps?

OpenStudy (anonymous):

I think I did something wrong...

OpenStudy (anonymous):

@iambatman @Hero

OpenStudy (anonymous):

\[f(x) = 8+\sqrt{5x-5}\] \[y = 8+\sqrt{5x-5}\] \[x = 8+\sqrt{5y-5}\] \[x - 8 = \sqrt{5y-5}\] \[(x-8)^2 = (\sqrt{5y-5})^2\] \[(x-8)^2 = 5y-5\] \[(x-8)^2+5 = 5y\] \[y = \frac{ (x-8)^2+5 }{ 5 }\] \[y = \frac{ 1 }{ 5 }(x-8)^2 \implies f ^{-1}(x) = \frac{ 1 }{ 5 }(x-8)^2\]

OpenStudy (anonymous):

@jtryon here are some steps that I used to achieve this. To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

yeah, I remember you helping another user with inverse functions so I was trying to follow the steps

OpenStudy (anonymous):

I gave you a step by step, so you can see how the process is, but next time you have to do this :P

OpenStudy (anonymous):

I did remember to replace f(x) with y and switch x with y in the function

OpenStudy (anonymous):

I think you got all that right, it was the algebra that gave you the trouble.

OpenStudy (anonymous):

Yeah, unfortunately I'm still rusty on the algebra and I really need to polish that up

OpenStudy (anonymous):

I have a bad habit of making my own algebra rules and breaking some

OpenStudy (anonymous):

Oh I see you switched the x's and y's but you were trying to simplify it or something, but your y disappeared lol from there, but in actuality you're suppose to solve for y.

OpenStudy (anonymous):

No worries :P, keep practicing!

OpenStudy (anonymous):

thanks :P

OpenStudy (anonymous):

I will use your example to study for my final

OpenStudy (anonymous):

Good luck friend! ^.^

OpenStudy (anonymous):

@iambatman, where did you get 1/5 for the last step in the algebra?

OpenStudy (anonymous):

Oh I see you found a mistake ;), nice job, at the end it should be \[f ^{-1}(x) = \frac{ 1 }{ 5 }((x-8)^2+5)\]

OpenStudy (anonymous):

That's your final answer.

OpenStudy (anonymous):

I'm still trying to figure out what step you did to get 1/5

OpenStudy (anonymous):

\[f ^{-1}(x) = \frac{ (x-8)^2+5 }{ 5 }\] You could just leave it as it is, I just factored the 1/5 out, you see dividing by 5 = 1/5.

OpenStudy (anonymous):

I'm good all the way up to that point

OpenStudy (anonymous):

Oh so you factored out a 1/5

OpenStudy (anonymous):

Yeah it doesn't matter though, the final answer I wrote is wrong, because I forgot to add the +5, but yeah you could leave it as I showed you just now.

OpenStudy (anonymous):

Hmm, it didn't take in that form so let me try with the 1/5 factored out

OpenStudy (anonymous):

Oh maybe you have to put it as \[f ^{-1}(x) = \frac{ 1 }{ 5 }(x^2-16x+69)\]

OpenStudy (anonymous):

OK, that is what they wanted

OpenStudy (anonymous):

I put it in with the 1/5((x-8)^2+5)

OpenStudy (anonymous):

Did that work? That's what happens when you're dealing with computers, it's never specific enough.

OpenStudy (anonymous):

Yeah, that answer with the 1/5 factored out did work.

OpenStudy (anonymous):

Alright cool, glad that worked :P

OpenStudy (anonymous):

Everything make sense now?

OpenStudy (anonymous):

I'll tag you if I get stuck on something else... have a few more problems to work through. It does so you were multiplying by the reciprocal of 5 which is 1/5